OpenStudy (anonymous):

A number decreased by 32 is -58. Find the number. The number is . (Give only the value of the number as your answer.)

OpenStudy (gschibby):

That one's easy ;) -58 + 32 = ?

OpenStudy (anonymous):

Just add -58+ 32

OpenStudy (anonymous):

-26

OpenStudy (gschibby):

Exactly! ^^

OpenStudy (anonymous):

thank you. I have another question, can you help?

OpenStudy (anonymous):

@Gschibby

OpenStudy (gschibby):

Sure!

OpenStudy (anonymous):

A blue bike is $14 less than a red bike. The sum of their prices is $300. How much is the red bike? The red bike is $. (Be sure to type only the price of the red bike. DO NOT use a dollar sign ($) in your answer I think it is 157?

OpenStudy (gschibby):

Okay, so the correct way to express this is (x = red bike, y = blue bike) x + y = 300 and x - y = 14 Solve the bottom one on y: y = x - 14 Insert that in the first one as y x + (x-14) = 300 solve it on x: 2x = 314 x = 314/2 x = 157 That is your x = red bike = 157 dollars add that to the second equation: 157 - y = 14 solve for y: y = 157 - 14 y = 143 dollars

OpenStudy (gschibby):

You can also solve this using matrixes, but this is a little less complicated ;)

OpenStudy (anonymous):

Thank you so much for your help :)